### v3

# Asymmetric Contrastive Multimodal Learning for Advancing Chemical Understanding

Abstract
The versatility of multimodal deep learning holds tremendous promise for advancing scientific research and practical applications. As this field continues to evolve, the collective power of cross-modal analysis promises to drive transformative innovations, leading us to new frontiers in chemical understanding and discovery. Hence, we introduce Asymmetric Contrastive Multimodal Learning (ACML) as a novel approach tailored for molecules, showcasing its potential to advance the field of chemistry. ACML harnesses the power of effective asymmetric contrastive learning to seamlessly transfer information from various chemical modalities to molecular graph representations. By combining pre-trained chemical unimodal encoders and a shallow-designed graph encoder, ACML facilitates the assimilation of coordinated chemical semantics from different modalities, leading to comprehensive representation learning with efficient training. We demonstrate the effectiveness of this framework through large-scale cross-modality retrieval and isomer discrimination tasks. Additionally, ACML enhances interpretability by revealing chemical semantics in graph presentations and bolsters the expressive power of graph neural networks, as evidenced by improved performance in molecular property prediction tasks from MoleculeNet. ACML exhibits its capability to revolutionize chemical research and applications, providing a deeper understanding of the chemical semantics of different modalities.
Keywords: Multimodal Learning, Molecular Graph Representation Learning, Chemical Data Mining, Contrastive Learning, Interpretability, Graph Neural Network.
1 Introduction
Multimodal deep learning (MMDL) is a thriving and interdisciplinary research field that is dedicated to enhancing artificial intelligence (AI) capabilities in comprehending, reasoning, and deducing valuable information across various communicative modalities [1]. It serves as an effective approach for the communication and integration of diverse data sources (text, images, audio, video, sensor data, etc), inspiring the innovative creation of more accurate and powerful AI models. For example, DALL-E2 [2] can showcase substantial progress in image generation and modification given a short text prompt. Witnessing the recent surge of research in image and video comprehension [3, 4], text-to-image generation [2, 5], and embodied autonomous agents [6, 7], multidisciplinary research, including biology, chemistry, and physics, also starts to embrace multimodal deep learning to unlock profound insights into complex systems, tackle challenging scientific problems, and push the boundaries of knowledge [8, 9, 10, 11, 12].
Figure 1: The framework of ACML (Asymmetric Contrastive Multimodal Learning). a. The conceptual view of ACML. Multimodal learning on multiple pairs of graph-chemical modalities (Smiles, Images, NMR, or Mass spectrometry) is established through asymmetric CLIP architecture [13]. The graph encoder is trained through ACML, while the encoder of chemical modality is pretrained and exempted from training. Embeddings from two modalities are projected into a joint latent space, aiming to align multiple views of the same molecule while distancing the embeddings from different molecules. b. Chemical knowledge transfer in ACML. For each chemical modality, an individual from-scratch graph encoder is paired and trained through the ACML framework, enabling it to express the learned chemical semantics from the corresponding chemical modality through latent embeddings. c. Comprehensive downstream tasks to demonstrate the effectiveness, interpretability, and generalization ability of graph learning in ACML.
In the realm of chemistry, MMDL has been extensively adapted to facilitate the representation learning across several chemical modalities, including chemical language, chemical notations, and molecular graphs [14, 15, 16, 17, 18, 19, 20]. It is essential to note that the potential of MMDL extends far beyond the above modalities [21, 22, 23, 24, 25, 26, 27, 28]. For instance, the molecular graphical depiction (referred to as image in the rest of the paper) [29] can offer powerful visual illustrations of molecules, providing unique insights into their spatial arrangements. Nuclear magnetic resonance (NMR) spectra, including 1H NMR and 13C NMR, offer a view of functional groups and characteristic features, enabling a deeper understanding of molecular behaviors and functionalities. Additionally, mass spectra from techniques such as gas chromatography–mass spectrometry (GCMS) and liquid chromatography–mass spectrometry (LCMS) provide a clue of molecular composition and fragmentation, aiding the identification and characterization of compounds. However, representing molecular information comprehensively using a single modality is a challenging task [30]. For example, string-based representations like SMILES [31] lack important topology information, whereas molecular depictions fail to include electronic information. MMDL could provide a solution by allowing the communication and integration of various heterogeneous modalities. However, there is limited work exploring such advantages in chemical systems, transcending the limitations associated with each modality.
Among the implementations of MMDL, contrastive learning has emerged as one of the most prevalent self-supervised approaches for information communication across multiple modalities [32]. Its objective is to align multiple views of the same instance across modalities while distancing the representations from different instances. One typical approach is to adopt a coordinated mechanism, like CLIP (Contrastive Language-Image Pretraining) framework [13], to enhance multimodal communication. The coordinated model establishes correspondences between the learned representations by first projecting these representations onto a joint hidden space and then contrasting positive pairs (similar instances) against negative pairs (dissimilar instances). In the domain of chemistry, great efforts have been directed towards using contrastive learning for tasks such as property predictions [16], unimodal pattern mining [33], and zero-shot classifications [21]. However, there remains a notable gap in research investigating the potential for integrating information across different modalities and elucidating the interpretability of learned representations within the context of multimodal coordination.
Many recent works have demonstrated strong advantages of molecular images [34], SMILES [35], GCMS [36], and molecular graphs [37] to produce molecular representations for drug discovery. But what contributes to such superiority? It’s important to recognize that the drug discovery process operates on multiple layers of informational hierarchy. These layers range from atomic-level properties such as hydrogen-bonding acceptors (HBA) to motif-level characteristics like hydrogen-bonding donors (HBD) and the number of stereogenic (chiral) centers, and finally to molecular-level attributes such as molecular weight, topological polar surface area (TPSA), and the logarithmic partition coefficient (LogP). To answer the question, we may need to investigate how diverse modalities impact on these hierarchical properties.
In light of these opportunities, we propose a novel approach called Asymmetric Contrastive Multimodal Learning (ACML) specifically designed for molecules. It leverages contrastive learning between the molecular graph and other prevalent chemical modalities, including G-SMILES, G-image, G-1H NMR, G-13C NMR, G-GCMS, and G-LCMS pairs (G for the abbreviation of Graph), to transfer the information from other chemical modalities into the graph representations in an asymmetric way. As graph representation can express hierarchical molecular information via carefully crafted graph neural networks (GNNs) [38, 39, 40, 41, 42, 43, 44, 45], we regard it as a valuable receptor containing basic topology information of a given molecule to assimilate information from other modalities. Unlike other initiatives in molecular properties prediction [46, 47], our emphasis is primarily on direct data mining utilizing graph latent representation. ACML enables graph representation learning to capture knowledge across various chemical modalities, promoting a more holistic understanding of the hierarchical molecular information within diverse input chemical modalities. This multimodal learning framework not only enhances the interpretability of learned representations but also holds the potential to significantly improve the expressive power of graph neural networks within the field of chemistry. Figure 1 illustrates the conceptual view of the ACML framework.
In our proposed framework, each ACML model involves the graph encoder and one chemical unimodal encoder, for example, G-SMILES represents the multimodal learning between graph and SMILES modalities. We utlize a dedicated pre-trained unimodal encoder for each chemical modalitiy (SMILES, Image, 1H NMR, 13C NMR, GCMS, and LCMS), keeping its parameters fixed in the ACML training stage. Concurrently, we train a simple and shallow graph encoder to effectively capture different representations of each modality. This approach allows the fixed chemical encoder to effectively transfer knowledge to graph encoders through asymmetric contrastive learning. Finally, we analyze and compare the embeddings generated by the trained graph encoders to discover their explanatory power. In summary, the advantages of the ACML framework can be viewed as follows:
  1. 1.
    Leverage effective asymmetric contrastive learning between graph and chemical modalities.
  2. 2.
    Achieve an efficient training scheme using the shallow graph encoder and pre-trained chemical encoders.
  3. 3.
    Empower knowledge transfer from chemical modalities into graph encoder during cross-modal chemical semantics learning.
  4. 4.
    Demonstrate the effectiveness and interpretability of graph learning in ACML, including three key tasks: (1) cross-modality retrieval (2) isomer discrimination, (3) revealing chemical semantics in graph embeddings, and (4) molecular property prediction.
2 Results
2.1 ACML framework
The proposed ACML framework is built upon the contrastive learning framework inspired by [13]. It effectively promotes correspondences between modalities of the same molecule (positive pairs) while contrasting them with different molecules (negative pairs). The whole pipeline, elaborated in Section 4.1, is composed of four components: an unimodal encoder for molecular modalities, a graph encoder formalized by graph neural networks, a non-linear projection head, and the temperature-scaled cross-entropy contrastive loss.
The ACML framework exhibits adaptability to diverse encoder designs, facilitating information extraction and compression. In our study, the encoders for each modality encompass Convolutional Neural Networks, trained to capture features for subsequent reconstruction tasks. Due to the inherent characteristics of each modality, the particulars of the encoder designs exhibit subtle variations. For instance, while 13C NMR and 1H NMR spectra both pertain to nuclear magnetic resonance, HNMR presents a higher degree of noise in comparison to 13C NMR, along with a greater prevalence of peak overlap. In contrast, 13C NMR peaks demonstrate enhanced distinctiveness. To address these disparities, 1H NMR data may require some preprocessing to extract peak-related, condensed information which leads to a lightweight encoder that strategically aims at extracting pivotal features from the peaks. For 13C NMR, the unaltered experimental data can be employed directly. The encoder architecture for 13C NMR has more convolutional layers, tailored to capture intricate details.
2.2 Cross-modality retrieval
Molecules can be represented by a variety of modalities, such as depiction images, SMILES notations, and traditional chemical characterization techniques. However, each modality possesses distinct capabilities and constraints when it comes to deciphering molecular structures. In the context of structural attributes, molecular depiction image and isomeric SMILES notation offer a detailed overview of the whole molecule from atom level to molecule level, facilitating the elucidation of their three-dimensional atomic configurations. NMR spectra provide distinct windows into the carbon and hydrogen atomic landscapes of the molecule, offering partial structural information about the molecule. Mass spectra provide crucial information about molecular mass and patterns of fragmentation, encompassing both molecular-level and motif-level information. To evaluate each ACML model, we conduct a cross-modality retrieval task: determine whether the model can accurately match a chemical modality to its corresponding graph modality from a large database of molecular graphs. Higher retrieval accuracy indicates better multi-modal coordination between graph and chemical representations. Note that the sample in the graph database never appears in the training phase.
Figure 2: Cross-modality retrieval accuracy for all ACML models, including G-SMILES, G-image, G-1H NMR, G-13C NMR, G-GCMS, and G-LCMS. In each ACML model. The X-axis represents the molecular pool ranging from 1000 (1K) to 1 million (1M) molecules. From this pool, the model is tasked to retrieve the target molecule by ranking the similarity scores. The Y-axis is the top k accuracy, which measures whether the target molecule is within the most similar k molecules the model recommended. All of our models can successfully identify the target molecule from more than 5000 candidates. Even when the molecular pool is as large as 1 million, models trained using the image and SMILES representations still achieve more than 90% accuracy.
Figure 2 presents the retrieval accuracy results for all ACML models, including G-SMILES, G-image, G-1H NMR, G-13C NMR, G-GCMS, and G-LCMS. These results are obtained by calculating the top k accuracy when matching one molecule’s chemical modality to its corresponding graph within a graph database of varying scales. The task becomes more challenging as the scale increases. When the scale transitions from 1 thousand (K) to 1 million (M), the G-Image gives the best results: the top 10 accuracy drops from 100% to 95.0%, which only experiences a reduction of 5.0%. G-SMILES gives the second best results, as top 10 accuracy drops from 100% to 87.5%. When it comes to G-1H NMR, the results are not as good as G-Image and G-SMILES, as expected. Specifically, when the scale transitions from 1K to 10K, the top 10 accuracy drops from 96.1% to 84.2%. And with 1M unseen molecules, the top 10 accuracy is 38.7%. G-13C NMR spectrum exhibits similar performance, as the top 10 accuracy from 96.3% to 83.9% until 43.4% when the scale transitions from 1K to 10K and up to 1M. Notably, G-13C NMR performs better than G-1H NMR when the search space expands. Intuitively, this aligns with the expectation, since the 13C NMR peaks are typically sharp and barely overlap in small molecules. On the contrary, the 1H NMR spectrums are more prone to signal overlap, especially in complex molecular structures. Although the encoders of 13C NMR and 1H NMR have been designed to cope with such differences, the training dataset inevitably falls short of encompassing the entirety of structural variations within the zero-shot validation set. Additionally, 1H NMR might encounter sensitivity issues, especially for compounds with low hydrogen content. Nonetheless, both the G-1H NMR and G-13C NMR exhibit satisfactory discrimination power in identifying the most probable graphs given the NMR sequences. This demonstrates the flexibility of our ACML framework with features from different spaces and dimensions generated from different designs of encoders.
When it comes to mass spectra modality, the results do not align as favorably as with images, smiles, and NMR spectra, as initially anticipated. For G-GCMS, when the scale transitions from 1K to 10k, the top 10 accuracy experiences a certain degree of reduction from 70.5% to 39.5%. When the validation database extends to 1M, the top 10 accuracy dropped to 7.8%. For G-LCMS, when the scale transitions from 1K to 10K, the top 10 accuracy experiences a certain degree of reduction from 32.4% to 14.9%. When the validation database extends to 1M, the top 10 accuracy dropped to 1.7%. One may observe that G-GCMS has better performance on molecular recognition than G-LCMS. The reason for that is GCMS is typically carried out under harsh conditions, resulting in a wide array of fragmentation that unveils the molecular composition. In contrast, LCMS is generally performed under milder conditions, resulting in reduced fragmentation. Particularly, LC-MS/MS primarily concentrates on localized information near the precise molecular weight. However, in terms of cross-modality retrieval, LCMS offers less comprehensive information.
2.3 Isomer discrimination
Besides calculating accuracy among a large group of molecules, in which many molecules could be intrinsically different, isomer recognition stands as one of the most challenging tasks in molecular identification due to the intricacy of isomeric structures. Distinguishing between stereoisomers, which share the same connectivity but differ in spatial arrangement, poses a particular challenge and demands advanced analytical methods and meticulous data interpretation. Case studies have been conducted to demonstrate the capabilities of NMR modalities in terms of isomer recognition. From the validation set, several pairs of complex isomers have been picked up randomly. These isomers not only have identical formulas but also feature highly similar functional groups. In this investigation, both 13C NMR and 1H NMR yield comparable results and effectively distinguish among the majority of instances. For each scenario, the ACML model is supplied with sets of NMRs, alongside confidence scores indicating the likelihood of these NMRs aligning with the corresponding pair of graphs. A greater confidence score signifies a stronger alignment according to our proposed framework. Four cases of 13C NMR are shown in Figure 3, where for each case, interested pairs of 13C NMR spectrum and possible molecular graphs are presented to the model. The model selects the molecule of the highest possibility to be the corresponding decipher of the 13C NMR spectrum. The proposed model correctly identifies 3 isomers out of 4. Similarly, four cases of 1H NMR are shown in Figure 4, where for each case, interested pairs of 1H NMR spectrum and possible molecular graphs are presented to the model. The proposed model correctly identifies 3 isomers out of 4.
Figure 3: 13C NMR Isomer discrimination task demonstrations. The correct pairs of 13C NMR and molecule are presented horizontally, linked by solid arrows. Out of the 4 challenging isomer pairs, our proposed model correctly identifies 3 isomers.
Figure 4: 1H NMR Isomer discrimination demonstrations. The correct pairs of 1H NMR and molecule are presented horizontally, linked by solid arrows. Out of the 4 challenging isomer pairs, our proposed model correctly identifies 3 isomers.
2.4 Visualization of molecular representations
In the ACML framework, the graph modality serves as an information receptor capable of bringing forth the information or chemical rules from another modality via coordination. Thus, it may reveal distinct dimensions of chemical significance among diverse chemical modalities. To illuminate variations of chemical semantics embedded in hidden graph representations acquired from various ACML frameworks, we computed graph embeddings of previously unseen 30,000 molecules using pre-trained graph encoders for analysis. Given 8 important chemical characteristics in drug discovery: molecular weight (MW), the logarithmic partition coefficient (LogP), hydrogen-bonding acceptor (HBA), hydrogen-bonding donor (HBD), topological polar surface area (TPSA), the number of rotatable bonds (#R-Bonds), and the number of stereogenic (chiral) centers (#C-Centers), we decided to examine how these characteristics correlate with the graph embeddings. The scope of this inquiry extends to understanding the impact of these modalities on crucial parameters within the domain of drug discovery, highlighting the strengths of each molecular modality.
We decomposed the feature dimensions via PCA [48] and mapped the resulting embeddings onto a two-dimensional plane. The outcomes were visualized as 2D points with coloring based on their corresponding chemical property values. Figure 5 presents the visualization results. Notably, a remarkable consistency in the distribution of points concerning the chemical property values was observed. This observation underscores the proficiency of the pre-trained graph encoder in capturing the inherent chemical semantics. We then conducted a quantitative evaluation of the relationship between graph embedding and the chemical properties using Pearson correlation coefficients (PCC) [49]. In our approach, we first established a linear regression model to determine the optimal linear combination of two PCA components that maximizes the Pearson correlation score between the linear combination of PCA components and chemical properties of interest. Table 1 presents the PCC scores of all ACML frameworks.
The knowledge of the chemical modalities was captured and highlighted by the degree of correlation, particularly with respect to key properties relevant to drug discovery. Regarding MW, G-Image achieved the highest score, followed closely by G-GCMS. Molecular weights can effectively be expressed by the pixels of molecular depiction image, and extracted from mass spectrometry data provided by GCMS. This accounts for the superior performance of these two models in this context. For HBA, G-SMILES emerged as the top performer. It is expected because electronegative HBA atoms, such as oxygen (O) or nitrogen (N), are directly denoted in SMILES. For HBD, G-Image demonstrates the best performances, as HBD groups, such as hydroxyl groups (-OH), amine groups (-NH2), and carboxylic acids (-COOH), are explicitly illustrated in the images. However, hydrogens are not always contained in SMILES, resulting in G-SMILES has a slightly lower performance. In the case of LogP, G-13C NMR and G-1H NMR are the game winners. This outcome is attributable to the fact that LogP measures the solubility of a solute in both water and organic solvents, making it highly reliant on electronic information within the molecule. This electronic information is well captured by the two NMR techniques, and it remains concealed in other modalities such as Image, SMILES, or mass spectrometry. For #R-Bonds, G-1H NMR exhibited the strongest performance because the rotation of chemical bonds can be derived from the magnetic equivalence of vicinal and geminal protons. When considering #C-Centers, G-SMILES achieves the highest correlation coefficient of 0.8297. This strong correlation can be elucidated by the inherent advantage of SMILES: atoms with chiral centers are directly annotated in SMILES representations, that is, “@" represents chirality with inward direction and “@@" represents chirality with outward direction.
Remark 1. Note that PCC only measures the linear correlation and doesn’t account for any nonlinear associations, a low PCC score doesn’t necessarily imply a lack of correlation. For example, when we examined the embedding visualization of G-1H NMR for the "#R-Bonds" task shown in Figure 5, we observed clear continuous color distributions, despite the PCC value being relatively low at 0.587.
Remark 2. These 8 chemical properties are very important in drug discovery. If the graph neural network is able to reveal this information, it offers the potential for deep models to understand the basic philosophy of molecular design and would even make breakthroughs in drug discovery. Note that these properties are never seen in ACML learning and the molecules we selected are also never seen in training, the surprising findings demonstrate our ACML framework can make a transition of chemical meaning to graph encoder, enhancing its interpretability.
Figure 5: Graph embedding visualization via PCA.
Table 1: Pearson correlation test results between graph representations and the chemical property of interest. The highest scores are marked as bold while the second highest are underlined. A higher score indicates that the low-dimensional graph representations have a higher linear correlation with the chemical property of interest.
MW #HBA #HBD LogP PSA #R-Bonds #C-Centers QED
G-Image 0.742 0.672 0.724 0.355 0.736 0.325 0.630 0.597
G-SMILES 0.623 0.892 0.716 0.564 0.869 0.244 0.830 0.492
G-1H NMR 0.202 0.361 0.227 0.588 0.303 0.587 0.388 0.235
G-13C NMR 0.195 0.673 0.569 0.805 0.659 0.277 0.641 0.188
G-GCMS 0.669 0.707 0.499 0.312 0.659 0.055 0.804 0.409
G-LCMS 0.615 0.254 0.120 0.338 0.229 0.087 0.769 0.292
2.5 Molecular property prediction tasks
Molecular property prediction is a critical graph-level prediction task used to assess the generalization capability of pre-trained graph encoders. In other words, it evaluates how well the encoder’s learned representations can be adapted to downstream tasks. In this section, we utilize the pretrained graph encoder from the ACML framework for downstream fine-tuning tasks, incorporating 6 classification and 2 regression benchmarks using MoleculeNet datasets [50]. For each task, we employ the shallow GIN [51] with 5 layers and 64 hidden dimensions as the backbone graph encoder, which is initialized with weights pre-trained by the ACML framework. We compare our models with other pretraining baselines on molecular graph learning in both classification and regression tasks. For classification tasks, ROC-AUC is used as the evaluation metric, and for regression tasks, RMSE is used. Among the baselines, L2P-GNN [52] and MGSSL [53] with different generation orders (BFS and DFS), MotifConv [44] propose self-supervised tasks within graphs to capture sub-structural or motif-level information during pretraining, while MolCLR [30] and MoMu [54] develop contrastive learning using augmented graphs or other modalities during pretraining. For a fair comparison, all baselines also use the 5-layer GIN model as the backbone graph encoder. We use scaffold splitting [55] to divide the dataset into the training set, validation set, and test set with a ratio of 8:1:1. For each dataset, the average and standard deviation of 5 individual runs are reported. More experimental details can be found in Section 5.2.2.
Table 2 shows the test performance on 8 molecular benchmarks. First, by comparing with no pretrained models, we observed that the best ACMLs generally outperform models without pretraining by big margins. Specifically, G-Image makes improvements in all datasets, including an average increase in AUC-ROC of 10.0% in BACE, 16.4% in Clintox, 7.9% in HIV, and reductions in RMSE with a 33.3% average decrease in ESOL and a 47.9% decrease in FreeSolv. These results suggest that contrastive learning with G-Image captures useful information in molecular depictions for a better understanding of the overall molecular structure. Second, by comparing with other self-supervised pretraining strategies, the proposed ACML framework achieves the best performance on 7 of 8 datasets. These improvements indicate that our ACML is a powerful self-supervised learning strategy, which is easy to implement and requires minimal manual design of self-supervised tasks. Third, by comparing ACMLs using different chemical modalities, we observed that G-Image generally outperforms other modalities as it achieves the best performance in BACE, BBBP, Sider, and ESOL, and ranks second in FreeSolv. In the Tox21 and HIV datasets, G-Image’s performance is not the highest, but it closely approaches the top results. This is likely due to its ability to capture comprehensive molecular structure information, which is crucial for most molecular property prediction tasks. These observations on the superiority of G-Image are also consistent with the analysis in Section 2.4: graph representations from G-Image exhibit high correlations with various chemical properties in drug discovery. In contrast, other modalities provide specific insights that could be beneficial for certain tasks but not universally effective. For instance, G-1H NMR achieves the highest average ROC-AUC of 83.9% in Clintox, 81.6% in Tox21, and 84.9% in HIV, but performs worse compared to the no-pratrained models in BBBP (3.3%) and Sider (3.3%). Similar cases are observed for other tasks and other modalities.
Table 2: Compare test performance (mean $\pm$ std) on molecular property prediction benchmarks. ROC-AUC (%) is for evaluations on BACE, BBBP, Clintox, Sider, Tox21, and HIV datasets. RMSE is used for evaluations on ESOL and FreeSolv datasets. The best result for each dataset is in bold. The second best is underlined.
BACE $\uparrow$ BBBP $\uparrow$ Clinton $\uparrow$ Sider $\uparrow$ Tox21 $\uparrow$ HIV $\uparrow$ Esol $\downarrow$ FreeSolv $\downarrow$
GIN (no pretrain) [51] 79.7 $\pm$ 4.9 87.3 $\pm$ 3.6 53.0 $\pm$ 6.5 61.6 $\pm$ 2.5 78.3 $\pm$ 2.4 76.2 $\pm$ 5.8 1.262 $\pm$ 0.156 4.672 $\pm$ 1.549
L2P-GNN [52] 81.9 $\pm$ 0.4 87.0 $\pm$ 0.8 54.0 $\pm$ 2.4 61.7 $\pm$ 1.8 77.4 $\pm$ 0.6 78.0 $\pm$ 2.6 1.772 $\pm$ 0.479 3.733 $\pm$ 1.323
MGSSL (DFS) [53] 79.7 $\pm$ 0.8 70.5 $\pm$ 1.1 79.7 $\pm$ 2.2 60.5 $\pm$ 0.7 76.4 $\pm$ 0.4 79.5 $\pm$ 1.1 0.914 $\pm$ 0.095 3.401 $\pm$ 0.896
MGSSL (BFS) [53] 79.1 $\pm$ 0.9 69.7 $\pm$ 0.1 80.7 $\pm$ 2.1 61.8 $\pm$ 0.8 76.5 $\pm$ 0.3 78.8 $\pm$ 1.2 0.861 $\pm$ 0.056 2.596 $\pm$ 0.997
MotifConv [44] 82.0 $\pm$ 5.5 90.0 $\pm$ 3.1 65.5 $\pm$ 13.9 62.7 $\pm$ 2.8 80.2 $\pm$ 1.5 80.0 $\pm$ 4.3 1.401 $\pm$ 0.153 3.697 $\pm$ 0.400
MoMu [54] 72.1 $\pm$ 0.3 70.2 $\pm$ 0.4 71.1 $\pm$ 3.0 56.1 $\pm$ 0.4 74.5 $\pm$ 0.5 77.2 $\pm$ 1.5 1.413 $\pm$ 0.104 3.374 $\pm$ 0.752
MolCLR [30] 74.6 $\pm$ 1.1 68.0 $\pm$ 2.0 62.0 $\pm$ 2.7 56.3 $\pm$ 2.2 73.9 $\pm$ 0.5 72.1 $\pm$ 1.3 1.354 $\pm$ 0.213 2.817 $\pm$ 0.911
G-Image 89.7 $\pm$ 0.8 87.4 $\pm$ 1.0 69.4 $\pm$ 2.3 62.9 $\pm$ 1.6 80.9 $\pm$ 0.8 84.1 $\pm$ 1.7 0.842 $\pm$ 0.002 2.431 $\pm$ 0.002
G-SMILES 85.6 $\pm$ 0.5 83.7 $\pm$ 0.5 68.5 $\pm$ 0.4 58.6 $\pm$ 0.4 81.0 $\pm$ 0.2 82.8 $\pm$ 0.7 0.940 $\pm$ 0.002 2.337 $\pm$ 0.002
G-1H NMR 86.6 $\pm$ 1.3 84.0 $\pm$ 3.8 83.9 $\pm$ 7.1 58.3 $\pm$ 0.9 81.6 $\pm$ 0.1 84.9 $\pm$ 0.9 1.059 $\pm$ 0.077 2.664 $\pm$ 0.301
G-13C NMR 87.5 $\pm$ 1.1 81.7 $\pm$ 3.2 75.6 $\pm$ 6.0 54.8 $\pm$ 0.8 81.6 $\pm$ 0.5 84.4 $\pm$ 0.7 1.049 $\pm$ 0.037 3.340 $\pm$ 0.314
G-GCMS 84.7 $\pm$ 2.1 79.1 $\pm$ 2.1 67.7 $\pm$ 8.9 60.3 $\pm$ 2.7 79.9 $\pm$ 1.1 83.7 $\pm$ 0.9 1.116 $\pm$ 0.038 3.807 $\pm$ 0.277
G-LCMS 79.7 $\pm$ 4.7 86.5 $\pm$ 3.4 62.3 $\pm$ 9.6 60.9 $\pm$ 0.6 81.4 $\pm$ 0.8 82.2 $\pm$ 0.5 1.108 $\pm$ 0.136 3.886 $\pm$ 0.196
3 Discussion and Conclusion
In this work, we introduced ACML framework, a novel approach tailored for molecular representational learning based on multi-modality contrastive learning. We demonstrated the inheritance of molecular graph-like structure allows graph representation to convey extensive information about molecules and treat the graph neural networks, i.e., the graph encoder, as a receptor to assimilate chemical semantics through multi-modal coordination, leading to effective and explainable molecular representational learning. Extensive experimental results on important chemical tasks, such as isomer discrimination, uncovering crucial chemical properties for drug discovery, and molecular properties prediction from MoleculeNet datasets, demonstrate the cross-modality transfer ability and the power of graph neural networks. In addition, our findings reveal a series of interesting questions to consider: of particular interest is how to effectively train a graph encoder to make it tailored to molecular representational learning. Instead of designing a deep and complex GNN framework and proposing various levels of self-supervised pretraining tasks, we demonstrate that the use of multi-modality contrastive learning can allow a shallow GNN with light training resources to produce an expressive graph encoder with great interpretability.
4 Methods
4.1 The ACML framework
We propose a Contrastive Multimodal Representation Learning (ACML) framework for molecules, which mainly contains three parts: (1) unimodal encoder of chemical modality; (2) graph encoder; (3) projection modules. The goal is to map the hidden representations of the molecular graph and the chemical modality into the joint latent space.
First, the inputs of each ACML framework consist of a molecular graph modality and one chemical modality. For each chemical modality, we make use of effective pre-trained encoders from publicly available sources, which have been demonstrated to be effective in the corresponding downstream tasks (See Table 3). As for molecular graph representation, we employ the graph neural network (GNN) as the graph encoder, where we elaborated several GNN variants in Section 4.2. These inputs are then fed into their respective encoders to produce embeddings. Mathematically, let $g$ and $c$ denote a molecule’s graph representation and its chemical modality representation, their embeddings after passing to their corresponding encoders are:
$\mathbf{h}_{g}=\text{ENCODER}_{\text{graph}}(g),\quad\mathbf{h}_{c}=\text{ENCODER}_{\text{chem}}(c)$ (1)
It’s worth noting that $\mathbf{h}_{g}\in\mathbb{R}^{d_{1}}$ and $\mathbf{h}_{c}\in\mathbb{R}^{d_{2}}$ may have different dimensions, i.e., $d_{1}\neq d_{2}$. To ensure that the information from two modalities could be coordinated in the same dimensional latent space, we apply two projection modules after both encoders. These modules assist in facilitating and coordinating between two modalities, resulting in the same dimensional hidden representations as projection outputs. In this study, the projection modules are designed as multiple-layer perceptrons (MLPs), which could be formulated as:
$\mathbf{h}_{g}^{\text{proj}}=\text{MLP}_{\text{graph}}(\mathbf{h}_{g}),\quad\mathbf{h}_{c}^{\text{proj}}=\text{MLP}_{\text{chem}}(\mathbf{h}_{c})$ (2)
where $\mathbf{h}_{g}^{\text{proj}}\in\mathbb{R}^{d}$ and $\mathbf{h}_{g}^{\text{proj}}\in\mathbb{R}^{d}$ have the same dimension.
Unimodal Representation Encoder Pre-trained Source
Image 2D image CNN Img2mol [29]
SMILES Sequence Transformer CReSS [56]
1H NMR Sequence 1D CNN N/Aa
13C NMR Sequence 1D CNN AutoEncoder [57]
GCMS & LCMS Sequence 1D CNN AutoEncoder [57]
aThe 1H NMR encoder undergoes pre-training using a CNN-based network similar to what’s described in [57]. Unlike the 13C NMR and GCMS/LCMS encoders, the 1H NMR encoder focuses on reducing the embedding dimensions to compress information while preserving satisfactory reconstruction ability. The diversity in encoder choices highlights the versatility of our suggested framework. This means encoders with varying degrees of modality information can be effectively integrated into the GNN network.
Table 3: Unimodal encoders of pervasive chemical modalities.
4.2 Graph encoder
Chemical molecules can be naturally represented as attributed relational graphs $G=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ and $\mathcal{E}$ are the node set and the edge set. Here a node $v\in\mathcal{V}$ and an edge $(v,u)\in\mathcal{E}$ represent the atom $v$ and the chemical bond connecting $u$ and $v$, respectively. The corresponding atom features (atomic number, chirality tag, hybridization, etc,.) and bond features (bond types, stereotypes, etc,.) are treated as node attributes and edge attributes in the graph. Graph neural networks (GNNs) [58, 59] which operate on graphs, have been combined with deep learning [60] to learn representations on graph data [61, 62, 63]. GNNs are also demonstrated as effective frameworks for molecular representational learning [64, 39].
Graph convolution. Many popular graph convolutional frameworks follow the message-passing scheme, including an iterative way of updating node representations based on the aggregations from neighboring nodes. Widely used approaches, such as GCN [65], GIN [51], GAT [66] and GraphSage [67], all follow message passing schemes. For the rest of the paper, we refer “message passing based GNNs” as “GNN”. For simplicity, we denote $\mathbf{x}_{u}^{(k)}$ as node feature vector of node $u$ in $k$-th layer, $\mathbf{e}_{v,u}$ as (optional) edge feature vector from node $v$ to node $u$, and let $\mathcal{N}_{u}$ represent the neighboring node set of the central node $u$. One message passing based convolution layer includes two steps: (1) create the message between the central node $u$ and its neighbor $v$, formalized by the function $\phi^{(k)}(\mathbf{x}_{u}^{(k)},\mathbf{x}_{v}^{(k)},\mathbf{e}_{v,u})$; (2) aggregate the messages of all neighbors of the central node $u$ using the operator $\bigoplus$, which denotes a differentiable and permutation invariant function, e.g., sum, mean or max operator. Mathematically, the latent node presentation is updated according to Equation 3 and will be fed into the next convolutional layer.
$\mathbf{x}_{u}^{(k+1)}=f^{(k)}(\mathbf{x}_{u}^{(k)},\bigoplus_{v\in\mathcal{N}(u)}\phi^{(k)}(\mathbf{x}_{u}^{(k)},\mathbf{x}_{v}^{(k)},\mathbf{e}_{v,u}))$ (3)
where $f^{(k)}$ and $\phi^{(k)}$ denote differentiable functions like MLPs (Multi Layer Perceptrons). $\mathbf{x}_{u}^{(k+1)}$ is able to captures the structural information within its $k$-hop network neighborhood after $k$ convolutional layers.
Graph pooling and readout. The node representations at the final iteration will be processed to generate the fixed-length graph-level representation $\mathbf{h}_{g}$, which involves graph pooling and readout operations. The pooling layer is usually applied to get a coarser graph which can be further reduced within a readout function, which is similar to conventional CNNs [68]. There is no strict distinction between pooling and readout operations in GNNs. Despite some studies [69, 70] that leveraged stacks of pooling layers for coarse graphs, most widely used works handling pooling and readout simultaneously by applying a permutation invariant function directly on all nodes in the graph to generate a fixed length graph representation [71, 72, 73, 74], which is generally formalized as
$\mathbf{h}_{g}=\text{READOUT}(\mathbf{x}_{u}^{(K)}\ |\ u\in\mathcal{V})$ (4)
The choice of aggregation operator. Many previous works [67, 51, 73, 75] demonstrate that the choice of aggregation function contributes significantly to the expressive power and performance of the model. In this work, we used sum operator in both message aggregation and graph readout function, which enables simple and effective learning of structural graph properties implied in [51].
4.3 Multimodal contrastive learning
Instance discrimination is a versatile technique that can be harnessed to contrast representations in various ways: within the same modality, across different modalities, or even through a combination of both modalities, referring to “joint" instance discrimination [76, 77]. The target of ACML is to learn a joint embedding between two different modalities while at the same time ensuring that similar features from the same modality stay close-by in the joint embedding. Therefore, in the implementation of the contrastive loss mechanism we adopt across-modality instance discrimination [78]. This intermodal contrastive strategy proves to be particularly beneficial in scenarios where understanding relationships and translating knowledge across modalities is a primary objective.
Mathematically, given a mini-batch of $N$ samples, $\{(g_{1},c_{1}),(g_{2},c_{2}),...,(g_{N},c_{N})\}$, where $g_{i}$ represents the graph modality and $c_{i}$ represents the chemical modality of the $i$-th molecule. According to Equ. 2, we denote $\{(\mathbf{h}^{\text{proj}}_{g_{1}},\mathbf{h}^{\text{proj}}_{c_{1}}),(\mathbf{h}^{\text{proj}}_{g_{2}},\mathbf{h}^{\text{proj}}_{c_{2}}),...,(\mathbf{h}^{\text{proj}}_{g_{N}},\mathbf{h}^{\text{proj}}_{c_{N}})\}$ as the hidden representation vectors after the projections. We denote the contrastive loss for the $i$-th graph representation as:
$L_{g_{i}}$ $=\sum_{j=1}^{N}-\frac{e^{s(g_{i},c_{j})/\tau}}{\displaystyle\sum_{k=1}^{N}e^{s(g_{i},c_{k})/\tau}}\cdot\log\frac{e^{\delta(g_{i},c_{j})/\tau}}{\displaystyle\sum_{k=1}^{N}e^{\delta(g_{i},c_{k})/\tau}}$ (5)
$=\sum_{j=1}^{N}-\text{softmax}(s(g_{i},c_{j})/\tau)\cdot\text{log}(\text{softmax}(\delta(g_{i},c_{j})/\tau))$
where $s(g_{i},c_{j})=((\mathbf{h}^{\text{proj}}_{g_{i}})^{T}\cdot\mathbf{h}^{\text{proj}}_{g_{j}}+(\mathbf{h}^{\text{proj}}_{c_{i}})^{T}\cdot\mathbf{h}^{\text{proj}}_{c_{j}})/2$, $\delta(g_{i},c_{j})=(\mathbf{h}^{\text{proj}}_{g_{i}})^{T}\cdot\mathbf{h}^{\text{proj}}_{c_{j}}$, and $\tau$ is the temperature parameter. Similarly, the contrastive loss for the $i$-th chemical modality representation is:
$L_{c_{i}}=\displaystyle\sum_{j=1}^{N}-\text{softmax}(s(g_{j},c_{i})/\tau)\cdot\text{log}(\text{softmax}(\delta(g_{j},c_{i})/\tau))$ (6)
Such loss design aims to guarantee that irrespective of the input source, whether it originates from two modalities, the optimization procedure is oriented towards attaining consistent representations.
5 Experimental details
5.1 Datasets
5.1.1 Training Datasets
For the molecular depiction image dataset, the molecules come from the Natural Products Magnetic Resonance Database (NPMRD) [79] and the images were produced by RDKit [80]. In total, there are about 270,000 samples in this dataset. For the SMILES dataset, molecules and SMILES come from NPMRD as well. In total, there are about 270,000 samples in this dataset. For 1H NMR dataset and 13C NMR dataset, the molecules, 1H NMR spectra, and 13C NMR spectra come from NPMRD as well. There are about 18,000 samples in 1H NMR dataset and 18,000 samples in 13C NMR dataset. For the GCMS dataset and LCMS dataset, molecules, LCMS spectra, and GCMS spectra come from MassBank of North America (MoNA) [81]. There are about 10,000 samples in the GCMS dataset and 13,000 samples in the LCMS dataset.
5.1.2 Zero-Shot Datasets
For the molecular candidate pool, 1M molecules were randomly chosen from PubChem, no intersection with the training dataset. For the spectra, 1000 spectra were used from the validation dataset.
5.2 Training Details
5.2.1 Pretraining in ACML
For graph modality, each atom is embedded by its atomic number and chirality type, and each bond is embedded by its bond type and stereotype. We used GNNs with ReLU activation function and a sum pooling is applied on each graph as the final readout operation to extract the hidden graph representation. The best model settings are provided in Table 4. Besides GIN [51], we tried multiple convolution techniques and found that GCN [65] achieved similar results to GIN, whereas GAT [66] and GraphSage [67] degraded performance during validation. We noticed that the number of layers did not have a strong effect as long as it exceeded a value of 5, implying that a shallow graph encoder is expressive enough. For chemical modalities, we used pre-trained unimodal encoders according to Table 3, where each encoder keeps frozen in the training phase. Outputs from the graph encoder and the chemical unimodal encoder are projected into the same dimensional joint hidden space using a 2-layer MLP with a specific projection dimension as provided in Table 4. All models were optimized with AdamW [82] in Pytorch implementation with $\beta_{1}=0.9,\beta_{2}=0.999$, and weight decay of 0.001. The model is trained with a batch size of 128 for a total of 100 epochs. Each model is trained on one NVIDIA Tesla V100 GPU using float32 precision. The entire framework is training efficiently since only one shallow graph encoder and two MLPs from the projection module need to be updated through training.
Model GNN type # Layers GNN hidden dim Projection dim
G-Image GIN 5 64 512
G-SMILES GIN 5 64 512
G-1H NMR GIN 3 512 256
G-13C NMR GIN 3 512 128
G-GCMS GIN 5 128 128
G-LCMS GIN 5 128 128
Table 4: Best Model settings in ACML.
5.2.2 Finetuning on molecular property prediction tasks
The following configurations were applied to all training tasks on the seven molecular benchmarks. We used a batch size of 32 and a maximal training epoch of 100. All experiments were conducted on one Tesla V100 GPU. We used the Adam optimizer. We tried different values of learning rate within [1e-2, 1e-3, 1e-4, 1e-5] and reported the best results. Before training, we performed data cleaning to remove certain molecules that failed to pass the sanitizing process in the RDKit or contained abnormal valence of a certain atom, as suggested in [83, 84, 44]. The detailed dataset statistics are summarized in Table 5.
Table 5: Summary of molecular property prediction datasets.
Dataset         Description # Graphs # Valid graphs # Tasks
BACE Quantitative (IC50) and qualitative (binary label) binding results for a set of inhibitors of human $\beta$-secretase 1 (BACE-1). 1513 1513 1
BBBP Binary labels of blood-brain barrier penetration (permeability). 2039 1953 1
Clintox Qualitative data of drugs approved by the FDA and those that have failed clinical trials for toxicity reasons. 1478 1469 2
Sider Database of marketed drugs and adverse drug reactions (ADR), grouped into 27 system organ classes. 1427 1295 27
Tox21 Qualitative toxicity measurements on 12 biological targets, including nuclear receptors and stress response pathways. 7831 7774 12
HIV Experimentally measured abilities to inhibit HIV replication. 41127 41125 1
Esol Water solubility data (log solubility in mols per litre) for common organic small molecules. 1128 1127 1
FreeSolv Experimental and calculated hydration free energy of small molecules in water. 642 639 1
6 Competing interests
All authors have no competing interests to declare.
7 Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used ChatGPT and Grammarly to grammar check and improve writing. After using this tool/service, the authors reviewed and edited the content as needed and took full responsibility for the content of the publication.
8 Data availability
The pre-training data from NPMRD and MoNA, and the zero-shot PubChem dataset are publicly available online, well-organized, and can be directly downloaded. RDKit can assist in parsing the dataset into individual samples.
9 Code availability
The code accompanying this work is available Github: https://github.com/GainGod-Xu/ACMLProject.